Suppose that the market basket for the university student price index (USPI) consists of 5 textbooks and 100 gallons of gasoline. In 2010, the base year for this index, textbooks cost $50 each and gas cost $1 per gallon. In 2011, textbooks cost $80 each and gasoline cost $3 per gallon. The USPI for 2011 is:
Answer #1

Cost of basket in 2010 = 5*50+100*1 = 350

Cost of basket in 2011 == 5*80+100*3= 700

CPI in 2011 = 700/350*100 = 200